OpenStudy (anonymous):

show that if abs(x+2)<2 then abs((x+2)(x+5))<4

OpenStudy (anonymous):

let t=x+2 so we have |t|<2 and we want to prove |t(t+3)|<4

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

I never tried it that way. I tried to show that since x+<1 yields x<-1 and that x+5<4 yields x<-1 and then I figured I needed to solve the expression (x+2)(x+5)<(x+5) x^2+7x+10<x+5 x^2+6x+5<0 (x+5)(x+1)<0 and since they both yield values less than 4 I figured that was true, but it was wrong

OpenStudy (anonymous):

emm...ignore my first reply :) lets do it with x

OpenStudy (anonymous):

Hahaha okie dokie then

OpenStudy (anonymous):

|x+2|<2 so -2<x+2<2 -4<x<0 right?

OpenStudy (anonymous):

Hmmm...so somehow that needs to translate to being <4?

OpenStudy (anonymous):

Wait, it is supposed to be \[\left| x+2 \right|<1\]

OpenStudy (anonymous):

Oops sorry for the confusion...haha

OpenStudy (anonymous):

now \[(x+2)(x+5)=x^2+7x+10=x^2+7x+\frac{49}{4}-\frac{49}{4}+10=(x+\frac{7}{2})^2-\frac{9}{4}\]and we have -4<x<0 so\[-\frac{1}{2}<x+\frac{7}{2}<\frac{7}{2}\]\[\frac{1}{4}<(x+\frac{7}{2})^2<\frac{49}{4}\]\[\frac{1}{4}-\frac{9}{4}<(x+\frac{7}{2})^2-\frac{9}{4}<\frac{49}{4}-\frac{9}{4}\]\[-2<(x+\frac{7}{2})^2-\frac{9}{4}<4\]so\[|(x+\frac{7}{2})^2-\frac{9}{4}|<4\]